# A Model Predictive Capture Point Control Framework for Robust Humanoid Balancing via Ankle, Hip, and Stepping Strategies

### v2

@@ VII-A3 Time horizon
VII-B Sequence for Ankle, Hip, and Stepping Strategies in Response to Disturbances
During robot walking, small CP errors induced by foot-ground impacts, model uncertainties, system noise, etc., can be effectively reduced through ankle strategy (ZMP control). However, large CP errors arising from external collisions or uneven terrain may not be adequately addressed by ankle strategy alone. In such cases, hip (CAM control) and stepping strategies (stepping control) are additionally employed. CP–MPC guides optimal behavior by leveraging the weights of the regulation terms in (10) ($\mathbf{w}_{\tau},\mathbf{w}_{F}$), determining which strategy to employ more actively.
In the hip strategy, after executing motions for balance recovery, there is a process of halting the robot’s motion and returning to the initial posture. Consequently, the hip strategy requires a certain amount of time to be reused. Moreover, excessive use of the hip strategy for small CP errors can lead to unnatural motions of the upper body. Therefore, we employ a variable weighting approach in Section IV-C to suppress the use of the hip strategy for small CP errors and allow its flexibility for large CP errors to facilitate balance recovery. As a result, depending on the magnitude of disturbances, the ankle and stepping strategies are employed initially, with the hip strategy used subsequently. This aligns with the preference for the stepping strategy following the ankle strategy in human behavior [6, 52, 53], as mentioned in Section VI-C.
VII-C CoM State Estimation
  • CoM position: The position of the CoM is estimated by utilizing global orientation information from the IMU and conducting forward kinematics calculations.
  • CoM velocity: The linear velocity of the pelvis is initially estimated using a complementary filter, integrating the linear acceleration from the IMU and the pelvis’s linear velocity that is kinematically calculated using joint angular velocities. The angular velocity of the pelvis is directly measured from the IMU. Subsequently, the velocity of the CoM is estimated by utilizing the spatial velocity of the pelvis, joint angular velocities, and the CoM Jacobian.
VII-D CAM Control using Upper Body Joints
In our framework, CAM control is performed using only upper body joints, and CAM generated from the lower body during walking is not explicitly considered. The CAM from the lower body is treated as a disturbance in CP–MPC and indirectly reflected in the CP error, which is typically compensated by ZMP control.
In the selection of upper body joints for CAM control, the yaw joint of the waist, neck joints, and wrist joints of both arms have been excluded, as they contribute minimally to CAM generation while introducing unnecessary computational load.
VII-E Adaptation to Early or Late Foot Contacts
@@
VIII Conclusion
This paper presents a novel balance control framework for humanoid robots to achieve robust balancing performance in the presence of disturbances. The proposed framework utilizes a linear MPC framework to drive the three balance strategies, namely ankle, hip, and stepping strategies, for CP tracking control (referred to as CP–MPC). This framework effectively calculates ZMP, CAM, and footstep positions to achieve CP tracking control. In addition, a novel variable weighting method that adjusts the MPC weighting parameters over the time horizon was proposed and applied to the CAM damping control. This method reduced the influence of damping action on CAM control and demonstrated better balancing performance against disturbances compared to the conventional constant weighting method.
Next, a hierarchical structure of CP–MPC and a stepping controller is proposed, enabling step time optimization based on CP–MPC variables. Specifically, the stepping controller adjusts the step time while closely following the footstep adjustments planned by CP–MPC to control the CP offset. The suitability of the parameter selection method used in the stepping controller is validated in comparison to the methods proposed in previous studies [26, 30, 31].
Finally, the balancing performance of the proposed method has been validated through simulations and real robot experiments under various scenarios with disturbances. The robot can traverse uneven terrain and overcome external forces. Furthermore, the superior balancing performance of the proposed method was experimentally verified, compared with a state-of-the-art QP-based CP controller that utilizes ankle, hip, and stepping strategies [39].
In the proposed framework, whole-body control is utilized for walking, but independent IK calculations are performed for the upper and lower body. Our future work includes implementing whole-body control based on inverse dynamics, rather than independent IK for the upper and lower body.
@@ VIII Conclusion
${\xi}_{T_{k},x}={A}^{T_{k}-k}\xi_{x,k}+\begin{bmatrix}{A}^{T_{k}-k-1}\mathbf{B}\cdots A^{0}\mathbf{B}\end{bmatrix}\begin{bmatrix}z_{x,k}\\ \tau_{y,k}\\ \vdots\\ z_{x,T_{k}-1}\\ \tau_{y,T_{k}-1}\end{bmatrix}$ (27)
@@
${p}_{ssp,x}=\frac{1}{1-{A}^{T_{k}-k}}\begin{bmatrix}{A}^{T_{k}-k-1}\mathbf{B}\cdots{A}^{0}\mathbf{B}\end{bmatrix}\begin{bmatrix}z_{x,k}\\ \tau_{y,k}\\ \vdots\\ z_{x,T_{k}-1}\\ \tau_{y,T_{k}-1}\end{bmatrix}.$ (29)

### v3

@@ VIII-A3 Time horizon
VIII-B Sequence for Ankle, Hip, and Stepping Strategies in Response to Disturbances
During robot walking, small CP errors caused by foot-ground impacts, model uncertainties, etc., can be reduced through ankle strategy (ZMP control). However, larger CP errors from external impacts or uneven terrain may exceed the ankle strategy’s capability, requiring additional use of hip (CAM control) and stepping strategies (stepping control). CP–MPC guides optimal behavior by leveraging the weights of the regulation terms in (10) ($\mathbf{w}_{\tau},\mathbf{w}_{F}$), determining which strategy to employ more actively.
In the hip strategy, after balance recovery motions, the robot halts and returns to its initial posture, requiring time to reuse. Moreover, excessive use of the hip strategy for small CP errors can lead to unnatural motions. Therefore, we employ a variable weighting approach to suppress the use of the hip strategy for small CP errors and allow its flexibility for large CP errors to facilitate balance recovery. As a result, the ankle and stepping strategies are employed first, with the hip strategy used as needed, similar to human behavior [6, 57, 58].
VIII-C CoM State Estimation
VIII-C1 CoM position
The position of the CoM is estimated by utilizing global orientation information from the IMU and conducting forward kinematics calculations.
VIII-C2 CoM velocity
The linear velocity of the pelvis is initially estimated using a complementary filter, integrating the linear acceleration from the IMU and the pelvis’s linear velocity that is kinematically calculated using joint angular velocities. The angular velocity of the pelvis is directly measured from the IMU. Subsequently, the velocity of the CoM is estimated by utilizing the spatial velocity of the pelvis, joint angular velocities, and the CoM Jacobian.
VIII-D Selection of IK over ID Control
Our research group is exploring high-performance whole-body ID control for humanoids using only proprioceptive sensors, without joint torque sensors, even in the presence of challenges such as friction and elasticity in high-reduction-ratio actuators. In this study, however, we opted for whole-body IK to develop the walking algorithm in parallel, without the presence of such uncertainty factors.
VIII-E Adaptation to Early or Late Foot Contacts
@@
VIII-F Limitations of balancing performance
From the external force experiments, we found that when addressing disturbances directed toward the stance foot, the kinematic constraints designed to prevent leg collisions hinder the adjustment of far footstep placements, degrading balancing performance. In such cases, strategies like rapid stepping or cross-leg movements could be beneficial.
In the uneven terrain experiments, our system, which employs position control with high gear ratio actuators, struggled to quickly respond to early contact with unexpected objects. As a result, unless the experimental walking duration is further extended, the success rate for overcoming objects over 3.5 cm diminishes. We plan to address this issue by researching whole-body ID control, as mentioned in Section VIII-D.
IX Conclusion
This paper presents a novel balance control framework for humanoid robots to achieve robust balancing performance. The framework employs a linear MPC to drive three balance strategies, namely ankle, hip, and stepping strategies, for CP tracking control, effectively calculating ZMP, CAM, and footstep positions. Additionally, a variable weighting method adjusts MPC weighting parameters over the time horizon for CAM damping control, reducing the influence of damping action on CAM and improving balancing performance.
Next, a hierarchical structure combining CP–MPC and a stepping controller is proposed, enabling step time optimization. Specifically, the stepping controller adjusts the step time while closely following the footstep adjustments planned by CP–MPC to control the CP offset. The parameter selection method in the stepping controller is validated against methods from previous studies [26, 30, 31].
Finally, the balancing performance of the proposed method has been validated through simulations and real robot experiments under various scenarios with disturbances. The robot can traverse uneven terrain and overcome external forces.
In the current walking framework, LIPFM dynamics are defined under the assumption of flat terrain, with no adjustments made for ground inclination affecting LIPFM dynamics, foot, or CoM trajectories. Our future work will adapt these dynamics and trajectories for sloped or uneven terrain. Additionally, to enhance generality, we plan to implement a state machine for step generation in response to disturbances, even after a predefined walking sequence or while standing still.
@@ IX Conclusion
${\xi}_{T_{k},x}={A}^{T_{k}-k}\xi_{x,k}+\begin{bmatrix}{A}^{T_{k}-k-1}\mathbf{B}\cdots A^{0}\mathbf{B}\end{bmatrix}\begin{bmatrix}z_{x,k}\\ \tau_{y,k}\\ \vdots\\ z_{x,T_{k}-1}\\ \tau_{y,T_{k}-1}\end{bmatrix}$